OpenStudy (anonymous):

Solve for x x^2(x-3)^2-4(x-3)^2=0

OpenStudy (anonymous):

\[x^2(x-3)^2-4(x-3)^2=0\] Factoring out (x - 3)², you have \[(x-3)^2(x^2-4)=0\]

OpenStudy (anonymous):

ok got that\

OpenStudy (anonymous):

so then you get (x^2-6x+9)x(x-4) right

OpenStudy (anonymous):

No, that doesn't help. What I said tells you that you have either \[(x-3)^2=0 \text{ or }x^2-4=0.\] Solving both for x gives you the solutions.

OpenStudy (anonymous):

I am stuck for the x(x-4) I got htat far what next

OpenStudy (anonymous):

Is it x=3 and x=-4

OpenStudy (anonymous):

Can someone please help me

OpenStudy (anonymous):

Sorry, was a bit busy. You're right about the x = 3 as a solution, but not with the other equation. \[x^2-4=0\\ x^2=4\\ x=\pm\sqrt{4}\\ x=-2,x=2\]

OpenStudy (anonymous):

so there are three answers x=-2 and x=2 and x=3

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

thanks